OpenStudy (anonymous):

simplify the logarithm log3(x+1) - log3(3x^2-3x-6)+ log3(x-2)

OpenStudy (anonymous):

PLEASE HELP!

OpenStudy (anonymous):

MEDAL TO ANYONE WHO ANSWERS

terenzreignz (terenzreignz):

Just remember these rules and you should be fine: \[\large \log_bM+\log_bN=\log_bMN\]\[\large\log_bM-\log_bN=\log_b\frac{M}{N}\]

OpenStudy (anonymous):

So I need to use both of those equations in order to get the answer?

terenzreignz (terenzreignz):

Well, use these 'properties'. So, in general, if it's a positive logarithm, it goes in the numerator, and if it's a negative logarithm, it goes in the denominator. This entire expression? It's actually very simple :) Just combine everything into one logarithm using those rules I posted...

OpenStudy (anonymous):

Okay! Thank you so much! I hope this works.

terenzreignz (terenzreignz):

You might want to stick around, though. At least tell me what you get when you combine them into one logarithm...

OpenStudy (anonymous):

Will do

OpenStudy (anonymous):

When setting it up should I have gotten Log3 (x+1)/(3x^2-3x-6) * log3 (x-2). ?

terenzreignz (terenzreignz):

That's right :) Let me... uhh... straighten that out for you :) \[\huge \log_3\frac{(x+1)(x-2)}{3x^2-3x-6}\]

terenzreignz (terenzreignz):

Now, were it as simple as this, the problem wouldn't be half as awesome :) Could you factor out that denominator, completely?

OpenStudy (anonymous):

No because you get a gcf of 3 and you're left with 3(x^2-x-2)

terenzreignz (terenzreignz):

By the way, you were only half-done with that... \[\huge \log_3\frac{x+1}{3x^2-3x+6}+\log_3{(x-2)}\] I hope you understand why I brought the (x-2) into the first logarithm. If you do, we'll proceed.

OpenStudy (anonymous):

Isn't it because the log bases are the same?

terenzreignz (terenzreignz):

Yeah, and the property, remember? \[\large \log_bM + \log_bN =\log_bMN\]

OpenStudy (anonymous):

Yep, got it!

OpenStudy (anonymous):

So now I have Log3 (x^2-2x+1x-2) / 3(x^2-x-2)

terenzreignz (terenzreignz):

So you said factor out the 3 in the denominator. Fair enough... \[\huge \log_3\frac{(x+1)(x-2)}{3(x^2-x-2)}\] But that's not done yet, you can still factor out that denominator.

OpenStudy (anonymous):

I have no idea haha.

OpenStudy (anonymous):

Wait no just kidding I do, it would be 3(x-2)(x+1) right,

terenzreignz (terenzreignz):

Hehe... oh you of little faith :) You're right, though, let's put that up... \[\huge \log_3\frac{(x+1)(x-2)}{3(x-2)(x+1)}\] And surely you must notice something :P

OpenStudy (anonymous):

THEY CANCEL OUT :D

terenzreignz (terenzreignz):

Yep, leaving you with...?

OpenStudy (anonymous):

Log3 3

terenzreignz (terenzreignz):

Whoops... a bit off, did you forget that 3 was in the denominator? :P

OpenStudy (anonymous):

Oh crap log3 (1/3) ?

terenzreignz (terenzreignz):

Very good. Now what is this equal to?

OpenStudy (anonymous):

I have no idea

terenzreignz (terenzreignz):

Well, what to what exponent do you raise 3 so that you get 1/3?

OpenStudy (anonymous):

-1 ?

terenzreignz (terenzreignz):

Bingo :P So that long complicated logarithm, was, in the end, just -1 The End XD

OpenStudy (anonymous):

Oh god bless. I hate math, you are amazing. Would you like to help me with more?

terenzreignz (terenzreignz):

Sure... hit me :D

OpenStudy (anonymous):

Psychologists use an exponential model of the learning process, f(t)= c(1-e^-kt), where c is the total number of tasks to be learned, k is the rate of learning, t is the time, and f(t) is the number of tasks learned. Suppose you move to a new school and you want to learn the names of 25 classmates in your homeroom. If your learning rate for new tasks is 20% per day, how many complete names will you know after 2 days? After 8 days?

terenzreignz (terenzreignz):

Sorry, Julia, I accidentally unplugged this laptop forgetting I removed the battery. I'm gonna need time to digest this...

OpenStudy (anonymous):

That's fine!

terenzreignz (terenzreignz):

Okay, it seems like your garden-variety plugging in task. You're given a function f(t) = c(1 - e^-kt) And you're to find the value of this function. Well, it's all in the power of interpretation! This would be a straightforward task if you know what c, k and t are, right?

OpenStudy (anonymous):

So you would solve it twice, once for 2 and 8?

terenzreignz (terenzreignz):

First of all, what are 2 and 8? Are they values for c? k? or t?

OpenStudy (anonymous):

T because it's time right?

terenzreignz (terenzreignz):

Very good :) So t = 2 and t = 8. What's c?

OpenStudy (anonymous):

25 I think

terenzreignz (terenzreignz):

I hope you're not guessing, but you're right. Your instincts must never fail you :) Now what's k, then?

OpenStudy (anonymous):

20% which would be .02

terenzreignz (terenzreignz):

No.... 20% is 0.2 You never said that ^ XD

OpenStudy (anonymous):

Oh whoops haha

OpenStudy (anonymous):

I got about 8 names

terenzreignz (terenzreignz):

Oh yeah, it won't be exact, since there's an e... anyway, I'll leave the computational bit to your calculator. I have faith in both of you :)

OpenStudy (anonymous):

Does 8 names and 20 names sound about right?

terenzreignz (terenzreignz):

I think so? I haven't input it in any calculator :)

OpenStudy (anonymous):

What is the inverse of log2 4x ?

terenzreignz (terenzreignz):

Okay, so you have y = log2 4x, right? To get the inverse, switch y and x x = log2 4y And try to isolate y :)

OpenStudy (anonymous):

Y= x-log2 /4 ?

terenzreignz (terenzreignz):

No... remember that the inverse of the logarithm involves exponents.

terenzreignz (terenzreignz):

Remember the definition of logarithm... \[\huge a=\log_bM \iff b^a=M\]

OpenStudy (anonymous):

In that case would a be y

terenzreignz (terenzreignz):

Well, to make it simple, make it so that a is x and M is 4y, and b = 2 And you'll get your answer.

OpenStudy (anonymous):

So y=2^x 4

terenzreignz (terenzreignz):

You mean \[\large \frac{2^x}{4}\] right? :P

OpenStudy (anonymous):

Oh.. That works too. Jesus I suck

OpenStudy (anonymous):

And would the inverse of y=log3 X be y=(1/3)^x ?

terenzreignz (terenzreignz):

Nope... try again :D

OpenStudy (anonymous):

Y=3^x ?

terenzreignz (terenzreignz):

Much better.

OpenStudy (anonymous):

The value of an industrial machine has a decay factor of 0.75 per year. After six years, the machine is worth $7500. What was the original value of the machine?

OpenStudy (anonymous):

I know you set it up with A=pe^rt but I'm not sure what to plug in since you're kind of working backwards with this one

terenzreignz (terenzreignz):

I don't know how to do this o.O

OpenStudy (anonymous):

Oh just kidding. What about (1/3)ln x + ln2- ln3 =3

terenzreignz (terenzreignz):

Properties of logarithms, just include this \[\huge p\log_bM=\log_bM^p\]

OpenStudy (anonymous):

I've never seen this before. Do you have a dumber version haha?

terenzreignz (terenzreignz):

This is as simple as it gets. In plain English, it says that if you multiply something to a logarithm, you can bring it up to become the exponent of that thing being, uhh, logaritmised :P

OpenStudy (anonymous):

So what would go where with the equation I have?

terenzreignz (terenzreignz):

With this... \[\huge \frac13\ln x=\ln x^{\frac13}=\ln \sqrt[3]{x}\]

OpenStudy (anonymous):

So is 3sqrt of x the answer?

terenzreignz (terenzreignz):

No. Far from it, you have to solve for x \[\huge \ln\sqrt[3]{x}+\ln2-\ln3=3\]

terenzreignz (terenzreignz):

Sorry, but I have to go to bed now... Us mere mortals need our sleep :) Good luck :) ---------------------- Terence out